please help me to this problem. express the density in g/cm3 of a 1.5kg object whose volume is 0.027 m3
Answer by mananth(16946)

express the density in g/cm3 of a 1.5kg object whose volume is 0.027 m3
-
Density = M / V
D= 1.5 /0.027
D= 1.5 * 1000 / 0.027 * 100 * 100 * 100
d=1500 / 27000
D= 55.55 kg/m^3
1 kg/m^3=0.001 g/cm^3
D=0.055g/cm^3
..